OpenStudy (nuccioreggie):

Solve the system of equations and choose the correct answer from the list of options. d + e = 15 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (10, −5) (5, 10) (10, 5)

OpenStudy (anonymous):

(10,5)

OpenStudy (sushi121212):

The answer would be (10,5). Plug in (10,5): d + e = 15 −d + e = −5 10+5=15 -10+5=-5 Which is correct because if you plug in all the other options, it wouldn't work.